OpenStudy (anonymous):

-4 (3 - p) > 5 (p + 1)

OpenStudy (anonymous):

show work also

OpenStudy (anonymous):

OK...you have on the LHS of the inequality: -4(3-p)=-4(3)-p(-4)=-12+4p For the RHS, you have 5(p+1)=5p+5

OpenStudy (anonymous):

LHS > RHS -12+4p > 5p+5 (subtract -4p from both sides) -12+4p-4p>5p-4p+5 -12>p+5 (subtract -5 from both sides) -12-5>p+5-5 -17>p

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

still confused??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

p is your variable...so the goal is to make the variable p on one side of the inequality...and other numbers on the other side

OpenStudy (anonymous):

oh ok